Good afternoon.
Here is a question I need some help with.
Find all of the zeros of F(x) if F(x)=(x^2-4)^3
This is what I have done so far
F(x)=(x^2-4)^3
F(2)=(2^2-4)^3
F(0)=(0^2-0)^3
Answer {2, (Triple zero), -2 (triple zero}
Any help is a lot of help so thank you much!

Answer by dabanfield(803)

Find all of the zeros of F(x) if F(x)=(x^2-4)^3

A zero of F(x) is a number k for which F(k) = 0.
We need to find values of x that cause (x^2-4)^3 = 0.
Note that x^2 - 4 = (x+2)*(x-2) so we need to find values of x such that
[(x+2)*(x-2)]^3 = 0
So zeros are where x-2 = 0 or x+2 = 0.
That is x = 2 or x = -2